A 1800 kg moving at 35 m/s. After 3.8 s the speed is 45 m/s. mat is the force experienced by B 4736.8 N C 3000.0 N D 3326.1 N A 4,655.2 N
Answer #1

Answer is B

by newton's first law of motion,

v = u + at

Here,

v= 45m/s

u= 35m/s

t= 3.8s

45 = 35 + 3.8a

a =10/3.8 m/s2

Force F = ma

        F = 1800   * 10 /3.8

         F= 4736.84N